### v1

# Shifting Attention to You: Personalized Brain-Inspired AI Models

Abstract
The integration of human and artificial intelligence represents a scientific opportunity to advance our understanding of information processing, as each system offers unique computational insights that can enhance and inform the other. The synthesis of human cognitive principles with artificial intelligence has the potential to produce more interpretable and functionally aligned computational models, while simultaneously providing a formal framework for investigating the neural mechanisms underlying perception, learning, and decision-making through systematic model comparisons and representational analyses. In this study, we introduce personalized brain-inspired modeling that integrates human behavioral embeddings and neural data to align with cognitive processes. We took a stepwise approach, fine-tuning the Contrastive Language–Image Pre-training (CLIP) model with large-scale behavioral decisions, group-level neural data, and finally, participant-level neural data within a broader framework that we have named CLIP-Human-Based Analysis (CLIP-HBA). We found that fine-tuning on behavioral data enhances its ability to predict human similarity judgments while indirectly aligning it with dynamic representations captured via magnetoencephalography (MEG). To further gain mechanistic insights into the temporal evolution of cognitive processes, we introduced a model specifically fine-tuned on millisecond-level MEG neural dynamics (CLIP-HBA-MEG). This model resulted in enhanced temporal alignment with human neural processing while still showing improvement on behavioral alignment. Finally, we trained individualized models on participant-specific neural data, effectively capturing individualized neural dynamics and highlighting the potential for personalized AI systems. These personalized systems have far-reaching implications for the fields of medicine, cognitive research, human-computer interfaces, and AI development.
Keywords Personalization  $\cdot$ Artificial Intelligence  $\cdot$ Cognition  $\cdot$ Transformers  $\cdot$ Magnetoencephalography
1 Introduction
Artificial intelligence (AI) systems are fundamentally composed of four key components: the model architecture that structures the processing mechanism, the learning rules that guide parameter updates, the objective functions that shape the learning path, and the datasets on which they are trained (Goodfellow et al., 2016; Saxe et al., 2021). While considerable progress has been made in each of these components and created much improved AI systems, aligning AI with human cognition remains a significant challenge (Ferrante et al., 2024; Lu et al., 2024; Rane et al., 2024). Such alignment holds the promise of not only enhancing the interpretability and robustness of AI models but also of deepening our understanding of human cognitive processes (Mahner et al., 2024).
Humans often rely on semantic abstraction and operate across varying levels of semantic categorization, making it essential for AI models to excel in both superordinate and fine-grained tasks (Goldstone, 1994; Schulte Im Walde and Frassinelli, 2022; Born et al., 2024). However, the representational strategies employed by humans and those utilized by AIs differ, highlighting the need for more sophisticated approaches to bridge this gap (Mahner et al., 2024). For instance, images that are indistinguishable to models–which form model metamers in later stages of neural networks–often appear obviously different to humans, demonstrating the presence of idiosyncratic, model-specific invariances that differ from the organization and representations of human sensory systems (Feather et al., 2023).
While the shared architectural principles between deep neural networks (DNNs) and human cortical processing (Cichy et al., 2016) play a role in brain alignment, recent studies suggest that the path to better human alignment may lie more in how models are trained than in their fundamental architecture (Muttenthaler et al., 2023a). The choice of training components, such as the objective function, significantly impacts the learning of human-like invariances (Nanda et al., 2023). Thus, one way to make models more human is to incorporate human inductive biases into AI systems to mirror cognitive processes. Indeed, leveraging these biases through Bayesian model selection and contrastive learning has been shown to enhance human alignment and machine learning performance (Sinz et al., 2019; Marjieh et al., 2023, 2024).
In this study, we aim to enhance the alignment between AIs and human cognition by fine-tuning the CLIP model (Radford et al., 2021) using human behavioral embeddings and neural data. Notably, the behavioral embeddings, derived from large-scale behavioral modeling (Hebart et al., 2019, 2023), provide interpretable dimensions that reflect human mental representations of objects. By training the model to incorporate this human-derived mental embedding space, we enable it to predict human behavior and neural representations more accurately, while also achieving significantly enhanced performance on alignment benchmarks (Fu et al., 2023).
While large-scale behavioral embeddings offer a fixed output of generalizable perceptual representations, neural data provides more dynamic insights that can reflect the millisecond-level temporal evolution of human brain representations (Ruffle et al., 2024). Moreover, neural data has the potential to better capture individual variability across diverse populations where behavior may be challenging to collect (e.g., infants, non-verbal population) (Kucyi et al., 2024a). To explore this, we further introduced a dynamic neural fine-tuning process using MEG signals (CLIP-HBA-MEG) to better align the model’s temporal representations with the rapidly changing activity patterns of the human brain.
Through training with group-level MEG representations (Hebart et al., 2023), we observed that the model generalizes well across external datasets, encompassing various participants, conceptual stimuli, and visual conditions. This finding reinforced our confidence in the robustness and generalizability of the approach, allowing us to move forward to individualized perceptual fine-tuning with participant-specific neural data and capture unique neural dynamics and perceptual differences. Our approach holds promise for creating personalized AI systems that reflect distinct and individual cognitive patterns, bridging the gap between artificial and biological cognition. By embedding these individualized patterns into AI models, we enhance their interpretability and alignment while deepening our understanding of the interplay between minds and machines.
2 Results
Figure 1: Schematic of the CLIP-HBA-Behavior fine-tuning process using behavioral data. The 66 SPoSE text dimensions are fed into the text encoder, producing 66 text representations $D_{1}\dots D_{66}$. Concurrently, visual stimuli from the THINGS dataset are input into the vision encoder, generating their corresponding visual representations $V$. These features from two modalities are bound via a dot product projection, mapping the visual features onto each of the text dimensions to form a 66-dimensional embedding $e$ for each image. Weight-Decomposed Low-Rank Adaptation (DoRA), a parameter efficient fine-tuning (PEFT) method, is used to fine-tune the attention modules of the text and vision encoder, using SPoSE behavioral embedding as an objective.
Current AI models trained on large-scale image datasets develop inductive biases based on the statistical patterns within those datasets. However, a significant gap exists between machine-derived inductive biases and the way humans develop theirs through perceptual experience (Goyal and Bengio, 2022). To address this, we introduced human-like inductive biases to the CLIP Model in the form of 66 Sparse Positive Similarity Embedding (SPoSE) dimensions, derived from large-scale behavioral studies (Zheng et al., 2019). These dimensions have been shown to effectively predict human similarity judgments for object distinctions (Hebart et al., 2020).
We developed CLIP-HBA-Behavior (Figure 1) by fine-tuning the CLIP model, using the 66 SPoSE dimensions as textual inputs while simultaneously feeding THINGS images (Hebart et al., 2019) as visual inputs. The CLIP architecture encodes both modalities—textual and visual—separately, before binding them via dot product similarity. This mechanism evaluates the encoded visual features against each of the 66 dimensions, resulting in a 66-dimensional embedding representation for each image input.
During fine-tuning, the model’s predicted embeddings are evaluated against the behavioral SPoSE embeddings using a Mean Squared Error (MSE) loss function. This loss guides the optimization process, updating the attention layers of both the text and visual encoders over multiple training iterations. The goal is to enable the model to accurately predict the SPoSE embedding for a given visual stimulus. In contrast to an off-the-shelf CLIP model, our behaviorally fine-tuned version, CLIP-HBA-Behavior, can produce significantly more human-aligned perceptual outputs. For example, when presented with an image of metal armor (Example Image Embedding in 1), our model identifies the object as highly associated with dimensions such as "metallic," "body-related," and "weapon-related" (top three dimensions), while correctly assigning low relevance to dimensions like "wood-related," "green," or "fluffy" (bottom three dimensions). See 8 for more examples.
2.1 Fine-Tuning CLIP with Human Behavioral Embeddings Enhances Alignment
Figure 2: Behavioral and neural alignment of fine-tuned CLIP-HBA-Behavior. (A) Behavioral results: Representational dissimilarity matrices (RDMs) for 48 objects, predicted by CLIP-HBA-Behavior (top) and CLIP-ViT (bottom), with Spearman rank correlations ($\rho$) of 0.78 and 0.32, respectively. (B) Neural results: Temporal correlations between model-predicted RDMs and MEG RDMs from THINGS (top left) and external datasets under varied conditions (remaining panels). CLIP-HBA-Behavior consistently outperforms CLIP-ViT in correlation and area under the curve (AUC), demonstrating enhanced neural alignment and generalizability.
As shown in Figure 2A, we evaluated the behaviorally fine-tuned model’s performance on a set of 48 object images from the THINGS database, which were fully sampled in a behavioral odd-one-out task (Hebart et al., 2020). These 48 objects were excluded from training to prevent data leakage. To assess the alignment between the model’s predictions and human behavior, we calculated the Spearman rank correlation ($\rho$) between the model’s predicted Representational Dissimilarity Matrix (RDM) (Kriegeskorte, 2008) and the ground-truth behavioral RDM. The behaviorally fine-tuned CLIP-HBA-Behavior achieved a Spearman correlation of 0.78 ($p<10^{-229}$) and a 95% confidence interval (CI) of [0.75, 0.80].
For comparison, we applied the same evaluation procedure to the original CLIP-ViT-L14 model. We extracted the 768-dimensional last-layer activations as the model’s best representation for each visual stimulus. The Spearman rank correlation between the CLIP-ViT model’s predicted RDM and the fully sampled behavioral RDM was 0.32, with a 95% CI of [0.26, 0.37]. This result highlights the significant improvement achieved by the CLIP-HBA-Behavior model, demonstrating over 100% improvement in generalizability and behavioral alignment compared to the baseline.
2.2 Improved Performance on Alignment Benchmark with CLIP-HBA-Behavior
Feature Type Model Name Val Score Test Score
768-d Visual Features CLIP-HBA-Behavior 0.88 0.88
CLIP-ViT-L/14 0.81 0.81
66-d SPoSE Dimensions CLIP-HBA-Behavior 0.85 0.84
CLIP-ViT-L/14 0.80 0.79
Table 1: Performance benchmarking on the NIGHT triplets dataset. The table compares validation (Val) and test scores for behaviorally fine-tuned CLIP-HBA-Behavior and the baseline CLIP-ViT-L/14 across two feature types: the 66-dimensional SPoSE semantic embedding space and the 768-dimensional visual features. CLIP-HBA-Behavior demonstrates a notable performance enhancement in generalizing to triplet-out decisions after behavioral fine-tuning.
To test our model’s robustness and performance, we benchmarked it against a state-of-the-art human visual similarity judgment metric—NIGHTS (Novel Image Generations with Human-Tested Similarity), a benchmark of 20,019 image triplets with human-tested perceptual similarity scores (Fu et al., 2023).
The results in Table 1 illustrate the significant improvements achieved by behaviorally fine-tuning CLIP-HBA-Behavior. To comprehensively evaluate the model, we tested two feature types. First, we assessed the 768-dimensional last layer activations of the Vision Transformer (ViT) encoder’s alignment with human perception. Second, we assessed the 66-dimensional SPoSE embedding space’s perceptual alignment. When using last layer visual features, CLIP-HBA-Behavior achieved validation and test scores of 0.88, surpassing the baseline score of 0.81. With the SPoSE dimensions, the model achieved validation and test scores of 0.85 and 0.84, compared to the baseline scores of 0.80 and 0.79. These results highlight that behavioral fine-tuning enhances human perceptual alignment and establishes CLIP-HBA-Behavior as a robust, generalizable model for nuanced perceptual distinctions.
2.3 Improved Neural Alignment with CLIP-HBA-Behavior
While the enhanced behavioral correlation of our model was anticipated due to its training on behavioral data, understanding whether the model’s embedding space truly becomes more human-like requires evaluating how training exclusively on behavioral data impacts its alignment with neural representations. Therefore, we further evaluated the alignment between the model’s representations against neural data obtained from MEG recordings (Hebart et al., 2023).
As shown in Figure 2B (top left), CLIP-HBA-Behavior demonstrated a stronger and more sustained temporal correlation with neural data across time points compared to the baseline CLIP model. Additionally, CLIP-HBA-Behavior’s correlation pattern and area under the curve (AUC) closely matched—and slightly exceeded—that of the SPoSE embedding model, whereas the CLIP-ViT baseline exhibited notably lower correlations and AUC.
To evaluate generalizability, we tested CLIP-HBA-Behavior on three external datasets containing different participants’ neural responses to out-of-distribution object stimuli under varying visual conditions:
  • Colored Images without Background (Figure 2B, top right): Clear object stimuli with backgrounds featuring humans, animals, fruits, and man-made objects (Grootswagers et al., 2017a).
  • Monochromatic Clear Images (Figure 2B, bottom left): Clear, monochromatic images of animate and inanimate objects (Grootswagers et al., 2017b).
  • Monochromatic Images with Blur (Figure 2B, bottom right): Blurry, monochromatic images of animate and inanimate objects (Grootswagers et al., 2017b).
In all cases, CLIP-HBA-Behavior outperformed baseline CLIP-ViT in terms of AUC. Notably, CLIP-HBA-Behavior consistently peaked in neural alignment around 300–400 ms after stimulus onset. This temporal pattern likely reflects the model’s fine-tuning on behavioral decisions and object semantics, which typically emerge later in processing, in contrast to earlier stages related to primary visual processing.
Overall, CLIP-HBA-Behavior demonstrated enhanced neural alignment across all datasets, spanning diverse image quality conditions and participant groups. Additionally, the performance benchmarking on the NIGHTS dataset highlights the significant improvements achieved by behavioral fine-tuning, with CLIP-HBA-Behavior outperforming the baseline model across both high-dimensional visual features and semantic embeddings. These results underscore that fine-tuning with human behavioral data not only improves the model’s ability to capture human cognitive representations but also establishes its robustness and performance in large-scale perceptual tasks. Additionally, we have found that behavioral fine-tuning using this pipeline can be effective with a behavioral dataset as small as 100 visual stimuli, making this method agile and suitable for specialized datasets of nuanced population groups that are likely more difficult to scale 10.
2.4 Dynamic Neural Fine-Tuning Captures Millisecond-Level Neural Dynamics
Figure 3: Schematic of the CLIP-HBA-MEG fine-tuning process using neural signals. A Feature Reweighting Matrix, pre-optimized at initialization, dynamically computes weighted combinations of vision encoder layer activations to align with neural decoding RDMs. Temporal scalers, $\alpha_{T}$ and $\beta_{T}$, respectively, modulate the magnitude of visual feature aggregation and the binding of visual-semantic features. Dimension-wise Gaussian noise is added to the embedding post-binding with a dynamically controlled noise level, mimicking varying stability of human neural responses while preventing model training from overfitting to specific noisy time points. Predicted RDMs are compared against MEG target RDMs via a custom loss function, updating the vision encoder’s attention layers and adapting the feature reweighting matrix. The text encoder is frozen with behaviorally pre-trained weights to ensure stable semantic features of text representations.
Building on the success of perceptual fine-tuning with behavioral data, which are large-scale, generalizable, and collected across many trials and participants, we sought to explore whether the model can improve by learning directly from human neural representations. Although typically collected from smaller participant samples, neural data can often be better suited to capture diverse neural profiles across different populations (Ye and Pandarinath, 2021). While behavioral embeddings provide static and generalizable perceptual representations, neural data offer dynamic insights, reflecting millisecond-level temporal patterns of human brain activity that cannot be captured through behavioral measures alone. Furthermore, neural data have the potential to better account for individual variability, providing a more personalized lens into how visual stimuli are processed (Pagan et al., 2024). To explore these possibilities, we developed a dynamic neural fine-tuning process using MEG signals, as demonstrated in Figure 3. The dynamically fine-tuned model, referred to as CLIP-HBA-MEG, adapts its visual stimulus embeddings to align with distinct stages of visual processing and the emergence of semantic representations.
2.5 Enhanced Behavioral and Neural Alignment after Dynamic Neural Fine-tuning
Figure 4: Behavioral and neural alignment of fine-tuned CLIP-HBA-MEG. (A) Behavioral validation: Comparison of the dynamic embedding space of the CLIP-HBA-MEG model across all timepoints (purple bars) with the THINGS behavioral data of 48 sample objects. The static behavioral alignment of the baseline CLIP-ViT model (green line) and the behaviorally fine-tuned CLIP-HBA-Behavior (red line) from Figure 2 are also plotted. CLIP-HBA-MEG demonstrates sustained higher behavioral alignment than the baseline soon after stimulus onset but peaks below the alignment achieved by CLIP-HBA-Behavior. (B) Neural validations: Temporal correlations between CLIP-HBA-MEG model-predicted RDMs and MEG RDMs are shown for THINGS images (top left panel) and external datasets under varied conditions: clear objects without background, clear monochromatic images, and blurry monochromatic images (remaining panels). CLIP-HBA-MEG consistently outperforms CLIP-ViT in correlation strength and area under the curve (AUC) across all tested conditions.
We evaluated the behavioral and neural validity of the CLIP-HBA-MEG model after fine-tuning with MEG responses from participants viewing visual stimuli in the THINGS image database. Figure 4A illustrates the temporal alignment of the dynamic embedding space of our model against the SPoSE behavioral embedding. Behavioral alignment increased post-stimulus onset, peaking around 600 milliseconds. The peak alignment ($\rho=0.65$) significantly surpasses the baseline CLIP-ViT’s static alignment ($\rho=0.32$, Figure 2A) but remains below the alignment achieved by the behaviorally fine-tuned CLIP-HBA model ($\rho=0.78$). These results suggest that, while training with neural data does not match the behavioral performance of the model compared to learning directly from behavioral data, it still significantly improves behavioral alignment over the baseline.
To assess the model’s robustness and generalizability in neural alignment, we tested it on the same three external visual datasets with varying characteristics and subjects used in Figure 2B.
As shown in Figure 4B, the neurally fine-tuned CLIP-HBA-MEG model with its dynamic embedding achieved higher overall neural alignment across all tested datasets, presenting higher peak alignment and overall AUC, compared to baseline CLIP-ViT’s static representation.
Overall, these findings confirm that the CLIP-HBA-MEG model not only maintains effectiveness in behavioral alignment, but also generalizes effectively across diverse visual conditions, aligning well with human perception and neural responses. However, we observed that the model’s neural alignment enhancement was less pronounced on degraded image datasets. This limitation likely stems from the neural fine-tuning process, which exclusively involved clear, naturalistic images from the THINGS database.
2.6 Viewable Dynamic Attention of CLIP-HBA-MEG
Figure 5: Example of dynamic saliency map of neurally fine-tuned CLIP-HBA-MEG
Leveraging the dynamic embedding spaces of the neurally aligned CLIP-HBA-MEG, we can visualize the specific regions and pixels the model prioritizes to produce human-like neural responses at a millisecond resolution. By combining randomized input sampling of the visual stimuli (Petsiuk et al., 2018) (Kaniuth et al., 2024a) with the model’s dynamically learned visual scaler, we can quantify both where and to what extent the model attends within any image following stimulus onset. As shown in Figure 5, the highlighted regions indicate where the model predicts human observers would direct their attention at a given time point.
2.7 Individualized Models Capture Participant-Specific Neural Dynamics
Figure 6: Individualized model schematic and results. (A) Fifteen individualized models were trained on MEG RDMs from 15 participants viewing 100 visual object stimuli. Each model achieved enhanced neural alignment specific to its corresponding participant, demonstrating sustained higher temporal alignment compared to the lower bound starting around 200ms post-stimulus onset. (B) The global correlation (RDMs) between the individualized models and their corresponding participant neural representations was assessed by comparing the differences among fine-tuned personalized models’ dynamic representations to the differences among the ground-truth temporal MEG signals of participants. Distances within the personalized models’ embedding spaces and participants’ MEG RDMs were computed by flattening along the time dimension and measuring Pearson distance between the resulting long vectors. This approach captures both time-specific representations and evolving patterns over time. When assesed using a set of 18 images excluded from training, the Spearman correlation between the RDMs of personalized models and participant neural RDMs was significantly high ($\rho=0.659,p<1\times 10^{-14}$), demonstrating that personalized neural fine-tuning successfully captured unique neural dynamics at an individual level.
We have demonstrated that our dynamic neural fine-tuning process, trained on group-level MEG data, generalizes well across multiple external datasets and diverse visual conditions, encompassing a wide range of participants and stimuli. These findings underscore the robustness and broad applicability of our approach, reinforcing confidence in its potential. Building on this, we sought to push the boundaries of perceptual fine-tuning by tailoring the model to individual neural profiles. By training on participant-specific MEG signals, we aim to capture unique differences in neural dynamics and signal processing at an individualized level, advancing toward our ultimate goal of creating personalized AI models that reflect the distinct cognitive representations and neural patterns of specific individuals.
As shown in Figure 6, we extended the dynamic neural fine-tuning pipeline of CLIP-HBA-Dynamic to train 15 personalized models, each fine-tuned on the individual MEG data of a single participant (Cichy et al., 2016). These personalized models achieved enhanced neural alignment specific to their respective participants, exhibiting sustained higher temporal alignment compared to the lower-bound noise ceiling, starting around 200 ms after stimulus onset.
To evaluate whether the personalized models captured consistent individual differences, we assessed the global correlation between the individualized models and the participants’ ground-truth neural representations. This was done by comparing the differences in dynamic embeddings of the fine-tuned models to the differences in temporal MEG signals among participants. Distances within the personalized models’ embedding spaces and participants’ MEG RDMs were calculated by flattening along the time dimension and computing the Pearson distance between the resulting vectors 11. This approach captured both time-specific neural representations and their evolving patterns over time. Validated on a set of 18 stimuli (excluded from training) viewed by the same 15 participants, the analysis revealed a significant Spearman correlation ($\rho=0.659,p<1\times 10^{-14}$) between the RDMs of the personalized models and the participant neural RDMs, confirming that individualized neural fine-tuning effectively captured unique neural dynamics and participant-specific patterns.
Figure 7: Individual Visual and Embedding Differences. A. Example of individual differences in dynamic attention: Personalized separate CLIP-HBA-MEG models fine-tuned on 2 individual subjects. B. Example of Temporal Embedding Individual Differences: Pearson correlation distance between 2 individual subjects’ model-predicted mental embeddings. C. Example of Time-Specific Individual Embedding Differences: 2 corresponding individual models for a given visual stimuli at a time point where their mental embeddings differ the most - 325ms.
With the individually aligned models, we can visualize nuanced differences in visual attention and dynamic mental embeddings at the subject level. By applying methods used to analyze model dynamic attention to each personalized model, we observe distinct visual dynamics across subjects following stimulus onset. As shown in Figure 7A, the personalized models of two different subjects exhibit varying temporal attention patterns while both successfully recognizing the presented objects.
Moreover, these personalized models enable the study of how subjects’ mental embeddings diverge and converge over time. In Figure 7, the two sample subjects’ mental embeddings begin to diverge around 100 milliseconds, peak at 325 milliseconds, and converge after 400 milliseconds. Figure 7C further illustrates dimensional differences in their mental embeddings. At the peak of divergence, Subject A’s representation is more skewed, relying on fewer dimensions, whereas dimensions such as "wood-related" rank higher in Subject B’s mental embedding. This suggests that at this specific timepoint for the example stimuli, Subject B is more likely integrating background context, while Subject A focuses primarily on the central object. Overall, these findings demonstrate that personalized models offer a powerful tool for studying cognitive processes at a millisecond-level resolution, providing insights into nuanced subject-specific variations.
3 Discussion
In this study, we have demonstrated that fine-tuning CLIP with human-derived behavioral and neural embeddings significantly enhances its alignment with human perception and opens the door for individualized AI models. CLIP-HBA-Behavior, trained on large-scale behavioral embeddings, showed a substantial improvement in predicting human similarity judgments while also indirectly capturing neural representational structure. When further refined with millisecond-level MEG data, CLIP-HBA-MEG achieved even greater alignment with human neural dynamics, successfully modeling the temporal evolution of perceptual representations. Notably, the model generalized well across external datasets with diverse participants, visual conditions, and object categories, reinforcing its robustness beyond the training set. Having established its capacity for human-aligned perception at the group level, we then adapted CLIP-HBA-MEG to capture individual differences in neural processing. By fine-tuning models on participant-specific neural data, we found that CLIP-HBA-MEG not only maintained strong alignment with group-level representations but also learned individualized neural dynamics, offering a promising path toward truly personalized AI systems.
3.1 Scalable and Transferable Applications
Building on these encouraging findings, our first area of focus was the scalability and transferability of the CLIP-HBA framework. CLIP-HBA-MEG leverages dynamic neural signals to capture real-time perceptual processes and higher-order conceptual organization, offering a nuanced window into human perception that static models cannot provide. In contrast to models that focus primarily on behavioral proxies—such as those that linearly map behavioral embeddings into model output (Kaniuth et al., 2024b) or perform generalizable local and global transforms to predict human similarity ratings (Muttenthaler et al., 2023b)—CLIP-HBA-MEG establishes a direct relationship with neural activity by explicitly capturing the emergence of information from dynamic neural signals. Recent research has shown that even when behavioral outcomes converge, the underlying neural dynamics can differ substantially between individuals (Renart and Machens, 2014; Bansal et al., 2021). Such findings underscore the critical importance of capturing dynamic neural activity. While earlier models—such as ReAlnet (Lu et al., 2024)—attempt to capture cross-modal EEG-fMRI alignments, they lack the temporal resolution necessary to track moment-to-moment perceptual shifts. Moreover, the CLIP-HBA framework can be fine-tuned sequentially, influencing the entire stream of information processing, which uniquely positions it as both a robust model of human perception and a versatile platform that can integrate diverse types of human behavior (e.g., MEG, MRI, behavioral data, electrophysiology). This adaptability is particularly useful in neuroscience applications, where neural data is often limited in scale.
3.2 Explainable Human-Driven AI Systems
In addition to capturing real-time perceptual processes, CLIP-HBA provides fully interpretable dimensions within the model embedding space (Hebart et al., 2020) that are absent in other models designed to incorporate human inductive biases (Lu et al., 2024). By integrating this embedding space with dynamic saliency mapping, we offer millisecond-level insights into both the model’s visual focus and the organization of its internal representations. Furthermore, findings indicate that when model decisions are transparent and systematically quantifiable, it becomes easier to evaluate the consistency of predictions, identify potential biases, and isolate errors for correction (Doshi-Velez and Kim, 2017; Rudin, 2019). These results support the view that interpretability is essential in applications where accurate, accountable decision-making is critical. Together, these advancements move AI closer to human-like cognition by promoting systems whose internal operations are more accessible and amenable to rigorous oversight (Zador et al., 2023; Sinz et al., 2019).
3.3 Expandable Architectures and Modalities for Future Directions
While our study demonstrates the effectiveness of fine-tuning CLIP to align AI models with human representations using behavioral and MEG data, our approach is not inherently limited to a specific model architecture, type of neural data, or sensory modality. The key factors contributing to the success of our method are the carefully designed loss functions and the modified mechanisms, such as feature reweighting (Kaniuth and Hebart, 2022) and noise injections. These techniques are broadly applicable to other transformer-based and deep neural network architectures, ensuring that our framework remains adaptable beyond CLIP. This allows our approach to be well-suited for expanding to additional sensory modalities. While our current work focuses on vision, future research could integrate other modalities such as auditory representations for multisensory learning (Baltrušaitis et al., 2018). By fine-tuning models with neural and behavioral data across multiple sensory domains, we can take the first steps toward a truly multisensory AI system—one that more closely mimics the way the human brain integrates information from diverse sensory inputs (Murray and Wallace, 2011; Tovar et al., 2020a). This expansion would not only enhance AI’s ability to model human perception but also open new possibilities for multisensory cognitive research and human-AI interaction.
3.4 Applications of Personalized Models
With the framework’s adaptability established, we now turn to its potential for personalization—a critical aspect of aligning AI with individual human cognition. The CLIP-HBA framework represents a significant advancement in personalized AI by integrating human behavioral embeddings with neural data to tailor models to individual cognitive profiles. This approach is deeply rooted in recent cognitive neuroscience research, which suggests that aligning artificial neural networks with biological cognition can enhance both interpretability and robustness (Doerig et al., 2023). By fine-tuning on millisecond-level MEG signals, the CLIP-HBA-MEG variant directly optimizes neural alignment, reflecting the neuroconnectionist framework’s emphasis on incorporating biological constraints rather than solely focusing on machine-centric performance. Moreover, the framework’s dynamic adaptation of learned representations based on neural feedback mirrors the functionality of digital twin systems (Jiang et al., 2021)—widely used in industrial and medical applications to model dynamic systems with real-time sensor data—thus acting as a cognitive digital twin that adjusts to an individual’s unique perceptual and neural idiosyncrasies.
Beyond the realm of neuroscience, personalized AI models like CLIP-HBA are making transformative strides in various domains, including education, healthcare, and industry. Advances in AI-driven learning style detection have demonstrated that machine learning can accurately classify users into distinct cognitive subtypes—such as visual, auditory, or kinesthetic learners (Kanchon et al., 2024)—which aligns with CLIP-HBA’s ability to fine-tune representations based on individual neural responses. Such personalization is critical for applications where individual differences matter; for example, in clinical contexts, these models can be harnessed to detect subtle deviations in neural dynamics associated with neuropsychiatric conditions (Damaraju et al., 2014; Preti et al., 2017), thus informing targeted interventions and enabling clinicians to monitor treatment efficacy over time. Furthermore, the capacity for real-time adaptation paves the way for advanced brain–computer interfaces that adjust to a user’s cognitive state, optimizing interaction and control in assistive technologies (Kucyi et al., 2024b). Collectively, these advancements underscore the transformative potential of personalized AI models to bridge the gap between artificial and biological intelligence, offering adaptive, interpretable, and human-aligned systems across a wide spectrum of fields.
3.5 Limitations and Need for Diverse Data
In thinking about the potential applications of personalized models, it is important to recognize the current limitations and the need for more diverse data. The success of our models in aligning with human cognition highlights the need for diverse and representative data. Incorporating behavioral and neural data from a broad participant pool reduces biases from homogeneous training sets, leading to more equitable and generalizable models (Born et al., 2024). However, beyond participant diversity, capturing the variability of real-world sensory experiences is equally important. As demonstrated in Figure 2B, CLIP-HBA-MEG, fine-tuned on group-level MEG data, exhibited strong alignment with human neural representations for clear, naturalistic stimuli but showed reduced generalizability on degraded visual datasets. This suggests that training exclusively on high-quality images may limit practicality and robustness. Since human perception naturally adapts to varied conditions—low lighting, occlusions, and noise—future datasets should reflect this variability by exposing the same participants to a range of visual conditions. This approach would enhance AI’s ecological validity, making models more adaptable and effective in real-world settings.
A major challenge in scaling these efforts lies within the difficulty of collecting large behavioral and neural datasets. While current datasets are often sufficient for traditional studies and comparative analyses, their limitations become increasingly prominent when trying to apply them to AI systems. To address this, exploring data augmentation techniques and leveraging generative AI to create synthetic neural and behavioral datasets could be a promising direction.
3.6 Conclusion
Our work underscores the versatility of the CLIP-HBA framework in adapting to a rich spectrum of data modalities, as demonstrated by its robust performance on large-scale, naturalistic stimulus collections. Looking ahead, systematic data collection efforts—ranging from active tasks within scanning environments to naturalistic interactions in immersive settings—hold the promise of revealing how human embeddings evolve when contextual demands change. By exploring active sensing, where representations directly support goal-directed actions, we may uncover entirely new dimensions of cognitive processing. The THINGS initiative (Hebart et al., 2019, 2023, 2020) has already paved the way for training models on diverse data types, and further endeavors that include neurodiverse populations and varied behavioral contexts will only enhance models like CLIP-HBA. Such efforts are not just a testament to the progress made so far but also a compelling challenge to further capture and understand the intricate dynamics of human perception and cognition.
4 Methods
4.1 Participants
Participants were drawn from existing datasets involving behavioral and neural measurements related to object recognition and mental representations. Behavioral data were obtained from a study that acquired extensive human similarity judgments of natural objects using the THINGS database. (Hebart et al., 2019, 2020). Neural data were sourced from magnetoencephalography (MEG) recordings collected in prior studies (Hebart et al., 2023) (Grootswagers et al., 2017a) (Grootswagers et al., 2017b) (Cichy et al., 2016). All participants provided informed consent in accordance with institutional guidelines approved by the relevant ethics committees.
4.2 Materials and Datasets
4.2.1 THINGS Dataset and SPoSE Behavioral Embeddings
The THINGS dataset consists of a comprehensive compilation of 1,854 natural object concepts accompanied by high-quality images and semantic embeddings derived from the Sparse Positive Similarity Embedding (SPoSE) model (Hebart et al., 2019). The SPoSE model deduces 66 interpretable human behavioral embeddings based on similarity judgments, capturing the semantic dimensions that structure human mental representations of objects. We subsampled 1806 object stimuli with their corresponding SPoSE embedding for the training and testing of CLIP-HBA-Behavior, excluding 48 fully-sampled stimuli data for measuring the model’s behavioral alignment.
4.2.2 MEG Data
We utilized two separate MEG datasets to fine-tune CLIP-HBA-MEG at both group and individual levels (Table 2). For group-level training, we subsampled averaged MEG RDMs from three participants and selected 1,806 out of 1,854 visual stimuli from the THINGS MEG dataset (Hebart et al., 2023), reserving the same 48 fully-sampled stimuli for behavioral validation. This dataset was chosen for its relatively large stimulus set, ensuring generalizability and providing a robust foundation for demonstrating the model’s ability to learn meaningful neural representations. For individual-level training, we used a dataset with 118 stimuli but a larger sample of 15 participants (Cichy et al., 2016), prioritizing the number of participants to assess the model’s capacity for individualized neural tuning and capturing inter-individual differences. To maintain consistency across training conditions and computational efficiency, we downsampled the individual-level dataset to 200Hz, matching the sampling rate used in group-level training.
Usage Group-Level Training Individual-Level Training
Source (Hebart et al., 2023) (Cichy et al., 2016)
# Participants 3 15
# Stimuli 1854 118
RDM Time Frame -100ms – 1300ms -100ms – 1000ms
RDM Sampling Rate 200Hz 1000Hz
RDM Distance Metric LDA SVM
Table 2: Summary of Neural Training MEG Datasets
Table 3 includes three external datasets, each with distinct participants, stimuli, and varying visual conditions. These datasets are used to evaluate the neural generalizability of CLIP-HBA-Behavior after behavioral fine-tuning and CLIP-HBA-MEG after group-level neural fine-tuning. By assessing performance on out-of-distribution data, they provide a critical test of the model’s ability to generalize beyond its training set. Both CLIP-HBA-Behavior and CLIP-HBA-MEG are designed to be robust to variations in the sampling rates of validation datasets, adapting dynamically through upsampling or downsampling within the sliding window mechanism to match the target dataset’s sampling rate.
Usage External Validation
Clear/No-Background Clear/Monochromatic Blurry/Monochromatic
Source (Grootswagers et al., 2017a) (Grootswagers et al., 2017b) (Grootswagers et al., 2017b)
# Participants 20 20 20
# Stimuli 36 48 48
RDM Time Frame -100ms – 600ms -100ms – 600ms -100ms – 600ms
RDM Sampling Rate 100Hz 100Hz 100Hz
RDM Distance Metric LDA LDA LDA
Table 3: Summary of External Neural Validation MEG Datasets
4.3 Model Architecture and Fine-Tuning Procedures
4.3.1 Baseline CLIP Model
We employed the Contrastive Language–Image Pre-training (CLIP) model as our base architecture (Radford et al., 2021). CLIP integrates a visual encoder, built on the Vision Transformer (ViT) architecture (Dosovitskiy et al., 2020), with a text transformer encoder. Pre-trained on a large dataset of image-text pairs, the model aligns visual and textual representations within a shared embedding space. For this study, we specifically employed CLIP-ViT-L/14, a variant with an increased parameter count to enable richer representational learning. This same model configuration serves as our baseline, ensuring a fair comparison when evaluating performance enhancements.
4.3.2 Fine-Tuning with Behavioral Data
Figure 1 illustrates a schematic of the fine-tuning process for the CLIP model, designed to enhance alignment with human behavior by leveraging SPoSE behavioral embeddings. The model processes inputs from two modalities—image and text—via separate streams. The 66 SPoSE dimensions are tokenized and passed through the text encoder, resulting in 66 textual representations, $D_{1}\dots D_{66}$. Simultaneously, image stimuli from the THINGS dataset are processed by the vision encoder, producing corresponding visual representations, $V$. The features from both modalities are then integrated through a dot-product projection, which computes the similarity between the visual features and text dimensions. This process encourages the model to process input stimuli into a human inductive bias mental embedding space. The feature binding function is defined as:
$e=\{V\cdot D_{i}\}_{i=1}^{66}$ (1)
where:
  • $\mathbf{e}$ is the model predicted 66-dimensional embedding vector for an input stimulus
  • $\mathbf{V}$ is the visual representation of the image stimulus, as produced by the vision encoder.
  • $\mathbf{D}=[D_{1},\dots,D_{66}]$ is the vector of 66 SPoSE dimensions, as produced by the text encoder.
The fine-tuning objective function uses Mean Square Error (MSE) loss, reinforcing the model’s predicted embeddings to closely match the SPoSE behavioral target embeddings. The MSE Loss is defined as:
$\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{e}^{(i)}-\mathbf{e}_{t}^{(i)}\right\|^{2}$ (2)
where:
  • $e^{(i)}$ is the model predicted 66-dimensional embedding for the input stimulus
  • $e_{t}^{(i)}$ is the target SPoSE 66-dimensional behavioral embedding for the same stimulus
  • $N$ is the total number of input stimuli in a training batch.
To balance fine-tuning effectiveness with computational efficiency, we employed Weight-Decomposed Low-Rank Adaptation (DoRA) (Liu et al., 2024) modules into the out-projection of the attention layers, specifically in the final text encoder layer and the last two vision encoder layers. DoRA extends the widely used parameter-efficient fine-tuning (PEFT) method, LoRA (Hu et al., 2021), by decomposing pre-trained weights into separate magnitude and directional matrices. This decomposition enables distinct magnitudinal and directional learning during fine-tuning, closely replicating the optimization dynamics of training directly on the original model weights. Only the DoRA parameters are updated, while the original layers remain frozen, significantly reducing the number of necessary training parameters and computational demands. Furthermore, to leverage CLIP’s pre-training on large image-text datasets, we limit the fine-tuning updates to only the last few layers of text and vision encoders, preserving transfer learning capabilities while maintaining the pre-trained model’s inherent generalizability (Fahes et al., 2024).
4.3.3 Behavioral Predictions Evaluation
We evaluated the fine-tuned CLIP-HBA-Behavior model’s ability to predict human behavior using a triplet odd-one-out task on a held-out portion of the data of 48 sample objects. (Hebart et al., 2020). This set was fully sampled to cover a wide range of naturalistic objects, and these 48 object images were specifically excluded during training to prevent any data leakage. Representational dissimilarity matrices (RDMs) were constructed from the model’s embeddings by calculating pairwise cosine distances (Kriegeskorte, 2008). These RDMs were compared to RDMs derived from actual behavioral choices using Spearman rank-order correlations (Zar, 1972), assessing the degree of alignment between model predictions and human judgments.
We evaluated the neural alignment of the fine-tuned CLIP-HBA-Behavior by computing the Spearman rank correlation between the RDM of the model’s static embedding space and the neural average-participant RDM representations at each timepoint. This analysis, conducted using the THINGS dataset and all external validation datasets listed in Table 3, assessed the effectiveness and generalizability of fine-tuning.
4.3.4 Dynamic Neural Fine-Tuning
We further developed a dynamic fine-tuning pipeline that enables the model to learn directly from neural representations and dynamics (CLIP-HBA-MEG) using a feature reweighting mechanism. This mechanism dynamically combines intermediate visual features from the vision encoder through a learnable feature reweighting matrix $\mathbf{W}$ Figure 3 illustrates this pipeline, where $\mathbf{W}$ operates across temporal and layer dimensions to align model representations with neural dynamics.
The feature reweighing matrix $\mathbf{W}\in\mathbb{R}^{T\times L}$ where $T$ is the number of MEG timepoints and $L=24$ is the number of visual layers in CLIP’s ViT-L/14 encoder. The matrix is initialized such that:
$W_{t,l}=\begin{cases}1&\text{if }l=L\\ 0&\text{otherwise}\end{cases}$ (3)
This initialization strategy assigns full weight to the final layer of the vision encoder while setting all other layers’ weights to zero, preserving a CLIP model’s pre-trained reliance on the last static representation. During training, $\mathbf{W}$ is parameterized as a neural network module with constraints enforced through architectural design: each row of $\mathbf{W}$ is normalized via min-max to ensure non-negative weights summing to 1.
Optimization follows a two-step gradient descent procedure designed to progressively align the model’s visual processing with human neural representations. Initially, only the feature reweighting matrix is optimized, shifting reliance from static last-layer representations to a dynamic aggregation of intermediate visual features. This enhances temporal dynamics while preserving the pre-trained capabilities of Vision Transformer (ViT) layer weights.
Stage 1 - Feature Reweighting: In this phase, the ViT parameters remain frozen while only the feature reweighting matrix $\mathbf{W}$ is optimized. To enforce temporal consistency, dependencies between neighboring time points are introduced through an average sliding window. Optimization is performed using the AdamW optimizer (Loshchilov and Hutter, 2019) for the initial epochs until convergence. This step establishes foundational temporal layer dependencies without altering the underlying visual processing capabilities of the pre-trained CLIP model.
Stage 2 - Joint Fine-Tuning: Once $\mathbf{W}$ is pre-optimized, all 24 ViT layers join the training process. Joint optimization is performed using DoRA with AdamW optimizer using a new set of learning rate, where $\mathbf{W}$ continues to be updated while the ViT layers undergo refinement. This approach enables the model to learn time-varying feature combinations through $\mathbf{W}$ while refining visual representations towards the target neural data. Detailed training hyperparameters (i.e. learning rates) can be found in Table 4.
The model incorporates temporal modulation through several key components: 1) visual feature magnitude scaling ($\alpha_{T}$) controlled by neural response richness (Carlson et al., 2011) (Carlson et al., 2013), and 2) semantic binding strength ($\beta_{T}$) determined by time generalization patterns (Tovar et al., 2020b) (King and Dehaene, 2014). Additionally, to better mimic the biological nature of human perception—where attention is never fully fixed on a stimulus while ignoring all else—we introduce a mechanism that allows the model to "think of something else" by incorporating dimension-specific Gaussian noise during semantic binding. This ensures a degree of randomness, resembling the natural variability in human neural processing:
$\mathbf{e}^{\prime}_{t}=\beta_{T}\cdot\left(\mathbf{e}_{t}+\boldsymbol{\epsilon}_{t}\odot\sigma_{\mathbf{e}_{t}}\right)$ (4)
where $\mathbf{e}_{t}$ represents the raw embeddings at time $t$, $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,1)$ is Gaussian noise, and $\sigma_{\mathbf{e}_{t}}$ denotes the standard deviation of embeddings across the batch dimension. The noise magnitude is dynamically scaled based on time-dependent neural stability, derived directly from target neural RDMs. This adaptive noise injection allows the model to develop data-driven variability while preventing overfitting to specific noisy time points, such as pre-stimulus recordings or moments long after stimulus onset. By incorporating this mechanism, the pipeline remains robust to variations in the timing of neural recordings, ensuring a more biologically plausible and flexible learning process.
MEG decoding RDMs serve as the target representations within this pipeline. Since neural data lacks a directly comparable 66-dimensional embedding space like the behavioral SPoSE embedding, we transform the model’s output embeddings into representational dissimilarity matrices (RDMs) using Pearson correlation (Kirch, 2008). This conversion translates the model embeddings into a common representational space with neural data, allowing alignment comparison against the target MEG decoding RDMs.
Given the dynamic nature of the model, its outputs and temporal MEG RDMs are both structured as a 3-dimensional object
$\text{RDM}_{\text{model}}^{\text{object}},\text{RDM}_{\text{MEG}}^{\text{object}}\in\mathbb{R}^{T\times N\times N}$
where:
  • $T$ is the number of timepoints
  • $N$ is the number of batch stimuli
This structure reflects the temporal evolution of representations. To construct reliable distance matrices for RDM analysis, a batch size greater than 1 is necessary. Maintaining a batch size of at least $N=32$ is optimal to ensure sufficient RDM sampling, which is critical for the stability of the training process.
A multi-objective loss function was incorporated to align the model’s representations with the temporal structure of the neural data. The loss function combines three complementary components: Pearson correlation loss, Mean Squared Error (MSE) loss, and Time Generalization loss.
The Pearson correlation loss ensures structural alignment between predicted and target representations while preserving temporal dynamics. For both model-predicted RDMs and target MEG RDMs, each RDM slice within the three-dimensional temporal RDMs is reduced to its upper triangle and vectorized into a one-dimensional vector (see Supplemental Figure 11). The Pearson loss is then computed as:
$\mathcal{L}_{\text{Pearson}}\left(x,y\right)=1-\frac{\sum_{i}\left(x_{i}-\bar{x}\right)\left(y_{i}-\bar{y}\right)}{\sqrt{\sum_{i}\left(x_{i}-\bar{x}\right)^{2}}\sqrt{\sum_{i}\left(y_{i}-\bar{y}\right)^{2}}},$ (5)
where $x$ represents the vectorized model-predicted RDMs, $y$ represents the corresponding vector of the target MEG RDMs, and $\bar{x}$ and $\bar{y}$ are their respective means.
The MSE loss complements the Pearson loss by regularizing the individual distances between stimuli within compared representations, defined as:
$\mathcal{L}_{\text{MSE}}\left(x,y\right)=\frac{1}{N}\sum_{i=1}^{N}\left(x_{i}-y_{i}\right)^{2},$ (6)
where $N$ is the total number of elements in the vectorized RDMs.
The Time Generalization loss regulates the temporal variation and stability of the model’s dynamic representation through a multi-step computation. First, we calculate a temporal representational similarity matrix (tRSM) by correlating each RDM slice with all other remaining timepoints’ slices:
$\text{tRSM}(t_{i},t_{j})=\frac{1}{N-1}\sum_{k=1}^{N}\left(\frac{\mathbf{r}_{t_{i}}^{(k)}-\mu_{t_{i}}}{\sigma_{t_{i}}}\right)\left(\frac{\mathbf{r}_{t_{j}}^{(k)}-\mu_{t_{j}}}{\sigma_{t_{j}}}\right)$ (7)
where $\mathbf{r}_{t}$ represents the vectorized upper triangle of the RDM at time $t$, $\mu_{t}$ and $\sigma_{t}$ are its mean and standard deviation, and $N$ is the number of stimuli pairs. The generalization profile $g_{t}$ is then obtained by averaging across the temporal dimension:
$g_{t}=\frac{1}{T}\sum_{j=1}^{T}\text{tRSM}(t,j)$ (8)
This captures how neural representations at each timepoint generalize across the temporal evolution. The final Time Generalization loss then measures the alignment between model-predicted ($g^{\text{pred}}$) and neural ($g^{\text{target}}$) profiles through:
$\mathcal{L}_{\text{Time Generalization}}=\mathcal{L}_{\text{Pearson}}\left(g_{t}^{\text{pred}},g_{t}^{\text{target}}\right)$ (9)
where $g_{t}^{\text{pred}}$ and $g_{t}^{\text{target}}$ are the respective generalization profiles of the model RDMs and the neural RDMs.
The total loss function for the MEG fine-tuning is:
$\mathcal{L}_{\text{Total}}=w_{1}\frac{\mathcal{L}_{\text{Pearson}}}{\mathcal{L}_{\text{Pearson}}^{0}}+w_{2}\frac{\mathcal{L}_{\text{MSE}}}{\mathcal{L}_{\text{MSE}}^{0}}+w_{3}\frac{\mathcal{L}_{\text{Time Generalization}}}{\mathcal{L}_{\text{Time Generalization}}^{0}},$ (10)
where $\mathcal{L}_{\text{Pearson}}^{0}$, $\mathcal{L}_{\text{MSE}}^{0}$, and $\mathcal{L}_{\text{Time Generalization}}^{0}$ represent the initial values of the Pearson loss, MSE loss, and Time Generalization loss, respectively. These values are used as normalization factors to ensure that each loss term is scaled consistently, preventing any single objective from dominating the overall optimization due to differences in magnitude.
$w_{1}$, $w_{2}$, and $w_{3}$ are hyperparameters that weight the contributions of the normalized losses. These parameters need to be adjusted based on different training datasets to ensure balanced optimization, allowing the three objectives to converge smoothly together.
4.3.5 Neural Data Alignment Evaluation
We validated the efficacy of our neurally fine-tuned CLIP-HBA-MEG model through both behavioral and neural alignment analyses. For the behavioral alignment evaluation, we used spearman correlation to compare every slice of the model’s dynamic RDMs against the static behavioral RDM derived the same 48 sampled stimuli, excluded from the neural fine-tuning process. For the neural alignment evaluation, we assessed the model’s performance by comparing its timepoint-to-timepoint RDMs with average-participant MEG decoding RDMs from the THINGS MEG dataset and all mentioned additional validation datasets mentioned in Table 3, using Spearman rank correlation (Zar, 1972).
4.4 Training Individualized Models
To account for individual differences in neural dynamics, we trained personalized models for each participant using the same CLIP-HBA-MEG pipeline used for group-level training, with a set of adjusted training hyperparameters 4 applied to all 15 individualize models. The key distinction was that, instead of using participant-averaged MEG decoding RDMs, each model was fine-tuned specifically on the MEG decoding RDMs corresponding to its respective participant’s neural data.
4.5 Evaluation of Individual Differences
Figure 6 presents the individual-by-individual dissimilarity matrices for both the neural data and the individualized models. For each pair of participants, we computed the Spearman correlation over time between the model-generated RDMs and the participant’s neural decoding RDMs used during training.
In analyzing the individualized models, temporal RDMs were flattened into one-dimensional vectors for both the model and target. We then computed Spearman correlations between individual models to construct a model-to-model dissimilarity matrix, using the Pearson correlation of the upper triangular elements of the temporal RDMs. Similarly, participant-to-participant RDMs were derived by first computing the Pearson correlation between objects, then flattening the timepoint and upper triangular RDM values into a single vector, as shown in Supplemental Figure 11. Both analyses offer a detailed comparison of representational similarity across participants and their respective models.
Finally, we assessed the correspondence between individualized model RDMs and participant RDMs using Spearman correlation, achieving a strong correlation of 0.65 when evaluated on a left-out validation set of 18 stimuli with the same participants’ neural data.
4.5.1 Dynamic Saliency Mapping of CLIP-HBA-MEG
We achieved millisecond-level visualization of dynamically important regions within any given visual stimulus using the fine-tuned CLIP-HBA-MEG model and Randomized Input Sampling for Explanation (RISE) (Petsiuk et al., 2018). The original RISE method identifies critical regions for model processing by comparing classification accuracy between unmasked images and randomly masked inputs.
Since CLIP-HBA-MEG functions as an embedding model without a classification head, we adapted this approach by measuring the cosine similarity between unmasked and masked outputs in the 66-dimensional embedding space. Specifically, if masking a region increases the cosine distance between the masked and unmasked embeddings, that region is inferred to be more crucial for the model’s ability to extract meaningful information and generate effective predicted embeddings. This masking process can be applied to any image at any timepoint within the model’s dynamic embedding space.
Additionally, we incorporated the learned visual scaler ($\alpha$) at each timepoint, allowing for a dynamically evolving visual strength modulated by target neural data. This integration enables us to capture the model’s "dynamic attention" as it processes visual information over time, as demonstrated in Figure 5.